OpenStudy (kcruz99):

A colony of bacteria is growing exponentially according to the function below, where t is in hours. About how many hours will it take for the population to reach 400? B(t)=4*e^0.8t A-12 B-18 C-24 D-6

OpenStudy (chrisdbest):

Looks like D is the closest

OpenStudy (nurali):

4*e^0.8t = 400 e^0.8t=100 ln (e ^ 0.8t) = ln (100) 0.8t * ln (e) = ln (100) 0.8t * 1 = ln (100) t = ln(100) / (0.8) = 1.25 * ln (100) t= 1.25 * 4.6 .: t=5.75≈6

OpenStudy (kcruz99):

Thank you so much!

OpenStudy (nurali):

My Pleasure.

OpenStudy (nurali):

Welcome to Openstudy.

OpenStudy (nurali):

Medal?

OpenStudy (chrisdbest):

Can I get a medal?

OpenStudy (chrisdbest):

@Nurali I need the medal more than you do.. lol